# Linearizability of flows by embeddings

### v6

@@ 3. Proofs
Next, assuming that Conditions 1 and 2 hold, we show that $(X,\Phi)$ is linearizable by a $C^{0}$ embedding. By Condition 2 and Theorem 2, there exists $n_{0}\in\mathbb{N}$, $B_{0}\in\mathbb{R}^{n_{0}\times n_{0}}$, and a $C^{0}$ embedding $F_{0}\colon A\to\mathbb{R}^{n_{0}}$ that linearizes $(A,\Phi|_{\mathbb{R}\times A})$. Since $X$ is homeomorphic to a subspace of a manifold, the Whitney embedding theorem implies the existence of a $C^{0}$ embedding $F_{1}\colon N\to\mathbb{R}^{n_{1}}$ with image $F_{1}(N)$ contained in the unit sphere $S^{n_{1}-1}\subset\mathbb{R}^{n_{1}}$ [Mun00, p. 316]. Thus, the map $F\colon X\to\mathbb{R}^{n_{0}}\times\mathbb{R}^{n_{1}}$ defined by
@@
$F\circ\Phi^{t}(x)$ $=(F_{0}\circ P\circ\Phi^{t}(x),e^{\tau\circ\Phi^{t}(x)}F_{1}(\Phi^{\tau\circ\Phi^{t}(x)}(x)))$
@@
for all $x\in A$ and $t\in\mathbb{R}$. This shows that the injective $C^{0}$ map $F$ is linearizing. We have shown that such an $F$ is proper, hence a $C^{0}$ embedding [Lee11, Cor. 4.97(b)].
Finally, assuming that $(X,\Phi)$ is linearizable by a $C^{0}$ embedding, we show that Conditions 1 and 2 hold. Restricting any linearizing $C^{0}$ embedding to $A$ and invoking Lemma 1 immediately yields Condition 2, so we need only verify Condition 1.
To do so, we may assume there is $n\in\mathbb{N}$ and $B\in\mathbb{R}^{n\times n}$ such that $\Phi$ is the restriction of the linear flow generated by $B$ to an invariant subset $X\subset\mathbb{R}^{n}$ that is locally compact in the subspace topology. Since $A$ is a globally asymptotically stable compact invariant set, $X$ must be contained in the sum of the real invariant linear subspace $E_{-}\subset\mathbb{R}^{n}$ corresponding to generalized eigenspaces of eigenvalues with negative real part for $B$ and the real invariant linear subspace $E_{0}$ corresponding to eigenspaces of purely imaginary eigenvalues. Moreover, $A\subset E_{0}$. By restricting $B$ to this sum, we may and do assume that $\mathbb{R}^{n}=E_{0}\oplus E_{-}$. Let
@@
As in the proof of Theorem 3, we may and do assume that all eigenvalues of $B$ have nonpositive real part. Let $E_{0}$, $E_{-}$, and
@@
be as in the proof of Theorem 3. The same argument from that proof implies that $P_{0}(X)=A$ and $P_{0}|_{A}=\textnormal{id}_{A}$, so that $P_{0}|_{X}$ is a continuous retraction when viewed as a map into $A$. But $P_{0}|_{X}$ is also $C^{k}$ when viewed as a map into $X$, which implies that $A$ is a $C^{k}$ embedded submanifold of $X$ [Hir94, p. 20]. Since $P_{0}$ commutes with the linear flow and hence $P_{0}|_{X}$ commutes with $\Phi$, Condition 1 holds with $C^{k}$ asymptotic phase map $P\coloneqq P_{0}|_{X}\colon X\to A$. Condition 1 and Theorem 1 in turn imply that Condition 2 holds.
@@
Next assume that Conditions 1, 2, 3 hold. Since the vector field generating $\Phi$ is uniquely integrable, there exists a proper strict $C^{k}$ Lyapunov function $V\colon X\to[0,\infty)$ for $A=V^{-1}(0)$ with respect to $\Phi$ ([Wil69, Thm 3.2], [FP19, §6], Remark 2). Since $V$ is proper, there is $c\in(0,\infty)$ such that $\{V\leq c\}\subset U$. A standard implicit function theorem argument implies that $N\coloneqq\{V=c\}\subset U$ is a $C^{k}$ embedded submanifold and the “time-to-impact-$N$” map $\tau\colon X\setminus A\to(-\infty,\infty)$ is $C^{k}$. Consider $F_{0}\colon X\to\mathbb{R}^{\ell}$ given by
@@
by Condition 3. Since $U$ and $X\setminus A$ are open, $F_{0}\colon X\to\mathbb{R}^{\ell}$ is $C^{k}$. Since $\tau(x)\to\infty$ and hence $F_{0}(x)\to\infty$ as $V(x)\to\infty$, $F_{0}$ is also a proper map.
By Condition 2 and Theorem 1, there exists $\ell_{1}\in\mathbb{N}$, $B_{1}\in\mathbb{R}^{\ell_{1}\times\ell_{1}}$, and a $C^{k}$ embedding $F_{1}\colon A\to\mathbb{R}^{\ell_{1}}$ that linearizes $(A,\Phi|_{\mathbb{R}\times A})$. Define the $C^{k}$ map $F\colon X\to\mathbb{R}^{\ell_{1}}\times\mathbb{R}^{\ell}$ by $F(x)\coloneqq(F_{1}\circ P(x),F_{0}(x)).$ Since $F_{0}$ is a proper map, so is $F$. Since
@@
Since $F|_{W_{0}}$ is an immersion and $F|_{A}$ is injective, there is an open set $W\subset A$ contained in $W_{0}$ such that the restriction $F|_{W}$ is injective. To show that $F$ itself is injective, fix $x,y\in X$ such that $F(x)=F(y)$. We will show that $x=y$ using injectivity of $F|_{W}$ and the fact that
@@
where we have used that $\tau\circ\Phi^{t}|_{X\setminus A}=\tau-t$. Global asymptotic stability of $A$ implies the existence of $s>0$ such that $\Phi^{s}(x),\Phi^{s}(y)\in W$. Since $F(x)=F(y)$ and (5) imply that $F|_{W}(\Phi^{s}(x))=F|_{W}(\Phi^{s}(y))$, injectivity of $F|_{W}$ and $\Phi^{s}$ implies that $x=y$. Thus, $F$ is injective. Similarly, taking tangent maps of both sides of (5) reveals that $F$ is an immersion. Since $F$ is also a proper map (as noted above), $F$ is a $C^{k}$ embedding [Lee11, Cor. 4.97] that linearizes $(X,\Phi)$ by (5). ∎
@@ 4. Conclusion
We obtained necessary and sufficient conditions for linearizability of $(X,\Phi)$ by a $C^{k\geq 0}$ embedding when $X$ is either compact or $\Phi$ has a compact global attractor, where $X$ is a $C^{k}$ manifold if $k\in\mathbb{N}_{\geq 1}\cup\{\infty\}$ and $X$ is a “reasonable” topological space if $k=0$ (Theorems 1, 2, 3, 4). We also obtained a separate sufficient condition for $C^{k\geq 1}$ linearizability in the compact case (Proposition 2), a necessary condition for the same (Proposition 1) having implications for linearizability in the presence of isolated equilibria (Corollary 1, 2, 3), a necessary condition for $C^{0}$ linearizability in the attractor basin case (Corollary 6), and extensions of the classical Hartman-Grobman and Floquet normal form theorems (Corollaries 4, 5). Additionally, we illustrated the theory in several examples (Examples 1, 2, 3, 4, 5, 6, 7, 8).
While linearizing embeddings have been studied for nearly a century, they remain a subject of interest, as explained in §1. In particular, in the field of applied Koopman operator theory, data-driven techniques involving the use of deep neural networks or algorithms like “extended Dynamic Mode Decomposition” (see, e.g., [LKB18], [BBKK22, §5.1, 5.4], [HK24]) are employed to compute linearizing embeddings. Importantly, however, such embeddings need not exist, and our results furnish precise conditions that determine when they do and do not. Failure for a dynamical system to satisfy these conditions imposes fundamental limitations on the mentioned data-driven techniques: they cannot possibly succeed at computing a finite-dimensional linearizing “embedding” of the dynamical system unless some of the usual requirements of an “embedding” (footnote 1) are relaxed, such as smoothness, continuity, or injectivity. These fundamental limitations complement other practical issues with applying such algorithms, such as those recently identified by Wu, Brunton, and Revzen [WBR21] and Haller and Kaszás [HK24].
Our results completely characterize linearizability by $C^{k\geq 0}$ embeddings for continuous-time dynamical systems having either a compact state space or at least one compact attractor (Theorems 1, 2, 3, 4, Corollary 6). While our results still furnish necessary conditions for the remaining case of a noncompact state space without any compact attractors, it would be interesting to completely characterize this remaining case as well. Additionally, even in the $C^{k\geq 1}$ global attractor case covered by Theorem 4, the relationship between one of our hypotheses and related literature remains to be fully understood (Remark 12).
@@ Acknowledgments
We thank Robert Cardona for communicating to us a generalized version of Tao’s universality notion as in §1, Zexiang Liu for useful discussions about the relation of our results with those of [LOS23a, LOS23b], and Paulo Tabuada for references to the control theory and polynomial flows literatures. We are also extremely grateful to Zexiang Liu for alerting us to an error in a previous version of Theorem 3.

### v7

@@ 3. Proofs
Next, assuming that Conditions 1 and 2 hold, we show that $(X,\Phi)$ is linearizable by a $C^{0}$ embedding. By Condition 2 and Theorem 2, there exists $n_{0}\in\mathbb{N}$, $B_{0}\in\mathbb{R}^{n_{0}\times n_{0}}$, and a $C^{0}$ embedding $F_{0}\colon A\to\mathbb{R}^{n_{0}}$ that linearizes $(A,\Phi|_{\mathbb{R}\times A})$. Since $X$ is homeomorphic to a subspace of a manifold, the Whitney embedding theorem implies the existence of a $C^{0}$ embedding $F_{1}\colon N\to\mathbb{R}^{n_{1}}$ with image $F_{1}(N)$ contained in the unit sphere $S^{n_{1}-1}\subset\mathbb{R}^{n_{1}}$ [54, p. 316]. Thus, the map $F\colon X\to\mathbb{R}^{n_{0}}\times\mathbb{R}^{n_{1}}$ defined by
@@
$F\circ\Phi^{t}(x)$ $=(F_{0}\circ P\circ\Phi^{t}(x),e^{\tau\circ\Phi^{t}(x)}F_{1}(\Phi^{\tau\circ\Phi^{t}(x)}\circ\Phi^{t}(x)))$
@@
for all $x\in A$ and $t\in\mathbb{R}$. This shows that the injective $C^{0}$ map $F$ is linearizing. We have shown that such an $F$ is proper, hence a $C^{0}$ embedding [1, Cor. 4.97(b)].
Finally, assuming that $(X,\Phi)$ is linearizable by a (necessarily proper) $C^{0}$ embedding, we show that Conditions 1 and 2 hold. Restricting any linearizing $C^{0}$ embedding to $A$ and invoking Lemma 1 immediately yields Condition 2, so we need only verify Condition 1.
To do so, we may assume there is $n\in\mathbb{N}$ and $B\in\mathbb{R}^{n\times n}$ such that $\Phi$ is the restriction of the linear flow generated by $B$ to a closed invariant subset $X\subset\mathbb{R}^{n}$. Since $A$ is a globally asymptotically stable compact invariant set, $X$ must be contained in the sum of the real invariant linear subspace $E_{-}\subset\mathbb{R}^{n}$ corresponding to generalized eigenspaces of eigenvalues with negative real part for $B$ and the real invariant linear subspace $E_{0}$ corresponding to eigenspaces of purely imaginary eigenvalues. Moreover, $A\subset E_{0}$. By restricting $B$ to this sum, we may and do assume that $\mathbb{R}^{n}=E_{0}\oplus E_{-}$. Let
@@
The proof that Conditions 1, 2, 3 in the following theorem are sufficient for linearizability by a $C^{k}$ embedding works by first constructing a linearizing $C^{k}$ embedding on a small neighborhood of $A$, then using a globalization technique of Lan and Mezić [27] (see also [28, 29, 55]) to construct a linearizing $C^{k}$ embedding defined on all of $X$.
@@
As in the proof of Theorem 3, we may and do assume that all eigenvalues of $B$ have nonpositive real part. Let $_{0}, EE_{-}$, and
@@
be as in the proof of Theorem 3. The same argument from that proof implies that $P_{0}(X)=A$ and $P_{0}|_{A}=\textnormal{id}_{A}$, so that $P_{0}|_{X}$ is a continuous retraction when viewed as a map into $A$. But $P_{0}|_{X}$ is also $C^{k}$ when viewed as a map into $X$, which implies that $A$ is a $C^{k}$ embedded submanifold of $X$ [2, p. 20]. Since $P_{0}$ commutes with the linear flow and hence $P_{0}|_{X}$ commutes with $\Phi$, Condition 1 holds with $C^{k}$ asymptotic phase map $P\coloneqq P_{0}|_{X}\colon X\to A$. Condition 1 and Theorem 1 in turn imply that Condition 2 holds.
@@
Next assume that Conditions 1, 2, 3 hold. Since the vector field generating $\Phi$ is uniquely integrable, there exists a proper strict $C^{k}$ Lyapunov function $V\colon X\to[0,\infty)$ for $A=V^{-1}(0)$ with respect to $\Phi$ ([69, Thm 3.2], [70, §6], Remark 3). Since $V$ is proper, there is $c\in(0,\infty)$ such that $V^{-1}([0,c])\subset U$. A standard implicit function theorem argument implies that $N\coloneqq V^{-1}(c)\subset U$ is a $C^{k}$ embedded submanifold and the “time-to-impact-NN” map $\tau\colon X\setminus A\to(-\infty,\infty)$ is $C^{k}$. Consider $F_{0}\colon X\to\mathbb{R}^{\ell}$ given by
@@
by Condition 3. Since $U$ and $X\setminus A$ are open, $F_{0}\colon X\to\mathbb{R}^{\ell}$ is $C^{k}$.
By Condition 2 and Theorem 1, there exists $\ell_{1}\in\mathbb{N}$, $B_{1}\in\mathbb{R}^{\ell_{1}\times\ell_{1}}$, and a $C^{k}$ embedding $F_{1}\colon A\to\mathbb{R}^{\ell_{1}}$ that linearizes $(A,\Phi|_{\mathbb{R}\times A})$. Define the $C^{k}$ map $F\colon X\to\mathbb{R}^{\ell_{1}}\times\mathbb{R}^{\ell}$ by $F(x)\coloneqq(F_{1}\circ P(x),F_{0}(x)).$ Since
@@
Since $F|_{W_{0}}$ is an immersion and $F|_{A}$ is injective, there is an open set $W\supset A$ contained in $W_{0}$ such that the restriction $F|_{W}$ is injective. To show that $F$ itself is injective, fix $x,y\in X$ such that $F(x)=F(y)$. We will show that $x=y$ using injectivity of $F|_{W}$ and the fact that
@@
where we have used that $\tau\circ\Phi^{t}|_{X\setminus A}=\tau-t$. Global asymptotic stability of $A$ implies the existence of $s>0$ such that $\Phi^{s}(x),\Phi^{s}(y)\in W$. Since $F(x)=F(y)$ and (5) imply that $F|_{W}(\Phi^{s}(x))=F|_{W}(\Phi^{s}(y))$, injectivity of $F|_{W}$ and $\Phi^{s}$ implies that $x=y$. Thus, $F$ is injective. Similarly, taking tangent maps of both sides of (5) reveals that $F$ is an immersion. Since $F$ is also a proper map by Theorem 3, $F$ is a $C^{k}$ embedding [1, Cor. 4.97] that linearizes $(X,\Phi)$ by (5). ∎
@@ 4. Conclusion
We obtained necessary and sufficient conditions for linearizability of $(X,\Phi)$ by a $C^{k\geq 0}$ embedding when $X$ is either compact or $\Phi$ has a compact global attractor, where $X$ is a $C^{k}$ manifold if $k\in\mathbb{N}_{\geq 1}\cup\{\infty\}$ and $X$ is a “reasonable” topological space if $k=0$ (Theorems 1, 2, 3, 4). We also obtained a separate sufficient condition for $C^{k\geq 1}$ linearizability in the compact case (Proposition 2), a necessary condition for the same (Proposition 1) having implications for linearizability in the presence of isolated equilibria (Corollary 1, 2, 3), a necessary condition for $C^{0}$ linearizability (Corollary 6), and extensions of the classical Hartman-Grobman and Floquet normal form theorems (Corollaries 4, 5). Additionally, we illustrated the theory in several examples (Examples 1, 2, 3, 4, 5, 6, 7, 8).
In particular, our results completely characterize linearizability by $C^{k\geq 0}$ embeddings for flows on connected state spaces that are either compact or contain at least one nonempty compact attractor (Theorems 1, 2, 3, 4, Corollary 6). While our results still furnish necessary conditions for the remaining case of a noncompact state space without any nonempty compact attractors, it would be interesting to fully characterize this remaining case as well. Additionally, even in the $C^{k\geq 1}$ global attractor case covered by Theorem 4, the relationship between one of our hypotheses and related literature remains to be fully understood (Remark 13).
Finally, we note that our results have implications for the field of “applied Koopman operator theory”, where algorithms like “extended Dynamic Mode Decomposition” ([71], [7, §5.1, 5.4], [48]) are used to attempt to numerically compute linearizing embeddings. Our results give precise conditions under which linearizing embeddings exist, imposing fundamental limitations on such algorithms complementary to practical limitations identified by practitioners [72, 48].
@@ Acknowledgments
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA9550-24-1-0299, managed by Dr. Frederick A. Leve. We thank Robert Cardona for communicating to us a generalized version of Tao’s universality notion as in §1, Zexiang Liu and Eduardo Sontag for useful discussions about the relation of our results with those of [40, 41], and Paulo Tabuada for references to the control theory and polynomial flows literatures. We are also extremely grateful to Zexiang Liu for alerting us to an error in a previous version of Theorem 3.